# Group class operations and homological conditions

### v1

@@ 2. A convenient class of groups for studying Gorenstein modules
We fix a commutative ring $k$ and let $\mathfrak{X}=\mathfrak{X}(k)$ be the class consisting of those groups $G$, for which ${\tt PGF}(kG)={\tt GProj}(kG)={\mathcal{P}}(kG)$. In other words, we ask that all cokernels of acyclic complexes of projective $kG$-modules be projectively coresolved Gorenstein flat. We also let $\mathfrak{X}^{\prime}=\mathfrak{X}^{\prime}(k)$ be the class consisting of those groups $G$, for which ${\tt GFlat}(kG)={\mathcal{F}}(kG)$ and $\mathfrak{Y}=\mathfrak{Y}(k)$ be the class consisting of those groups $G$, for which ${\tt GInj}(kG)={\mathcal{I}}(kG)$. In other words, $G$ is an $\mathfrak{X}^{\prime}$-group (resp. a $\mathfrak{Y}$-group) if all cokernels (resp. kernels) of acyclic complexes of flat (resp. injective) $kG$-modules are necessarily Gorenstein flat (resp. Gorenstein injective). These classes have been studied in [22], where the following result is proved.
@@
(i) $\mathfrak{X}=\mathfrak{X}^{\prime}$; cf. [22, Corollary 2.3].
(ii) $\mathfrak{X}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$, $\Phi$ and $\Phi_{flat}$; cf. [22, Theorem 3.3]
(iii) $\mathfrak{Y}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi_{inj}$; cf. [22, Theorem 4.1].
For later use, we record the following properties of Gorenstein projective modules over a group ring $kG$. The first three of these properties are valid over all groups, whereas the latter three are valid under additional assumptions:
@@
(5) If $G$ is an $\mathfrak{X}$-group, then $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is a complete hereditary cotorsion pair.
@@
Proof. Property (1) is precisely [21, Lemma 2.6(i)], whereas (2) and (3) are immediate consequences of the definition of Gorenstein projectivity; see also [27, Theorem 2.20]. Property (4) follows since restriction maps $\mathcal{P}(kG)$ into $\mathcal{P}(kH)={\tt GProj}(kH)$, whereas (5) and (6) follow since ${\tt GProj}(kG)={\tt PGF}(kG)$, in view of [35, Theorems 4.4 and 4.9]. $\Box$
@@
(5’) The pair $\left({}^{\perp}{\tt GInj}(kG),{\tt GInj}(kG)\right)$ is a complete hereditary cotorsion pair.
Proof. Property (1’) follows by dualizing the argument in [21, Lemma 2.6(i)], whereas (2) is an immediate consequence of the definition of Gorenstein injective modules. Properties (3’) and (5’) follow from [35, Theorem 5.6] and [27, Theorem 2.22]. Finally, property (4’) follows since restriction maps $\mathcal{I}(kG)$ into $\mathcal{I}(kH)={\tt GInj}(kH)$. $\Box$
@@ 3. $\mbox{Ext}^{1}$-orthogonality and group class operations
Let $\mathfrak{X}_{0}$ be a subclass of $\mathfrak{X}$. We consider a group $G$ and a $kG$-module $M$, such that $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
@@
(iv) We have $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\Phi_{flat}\mathfrak{X}$-subgroup $H\subseteq G$.
@@
for such a subgroup $F$ and all $n>0$. It follows that $\mbox{Ext}_{kH}^{n}\!\left(N_{i},\mbox{res}_{H}^{G}M\right)\!=0$ for all $i$ and all $n>0$. Then, using dimension shifting, we conclude that $\mbox{Ext}_{kH}^{n}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>d$. Since this holds for any $N\in{\tt GProj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${\tt GProj}(kH)^{\perp}$. Indeed, the Gorenstein projective $kH$-module $N$ is the $d$-th syzygy in a projective resolution of another Gorenstein projective $kH$-module $N^{\prime}$ (this is property (2) in Section 2) and hence
@@
Here, $K$ is equal to the continuous ascending union of its submodules $(K_{\lambda})_{\lambda<\kappa}$. Since $K_{\lambda+1}/K_{\lambda}$ can be identified with the kernel of the (surjective) structure map $N_{\lambda}\longrightarrow N_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kH}(K_{\lambda+1}/K_{\lambda},\mbox{res}_{H}^{G}M)=0$ for all $\lambda$ and $n>0$. Therefore, Eklof’s lemma [24, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kH}\!\left(K,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>0$. Then, the short exact sequence (1) above implies that $\mbox{Ext}^{n}_{kH}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>1$. Since this holds for any $N\in{\tt GProj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${\tt GProj}(kH)^{\perp}$. Indeed, the Gorenstein projective $kH$-module $N$ is the first syzygy in a projective resolution of another Gorenstein projective $kH$-module $N^{\prime}$ (this is property (2) in Section 2) and hence
@@
We shall prove in the next Section (Proposition 4.3) that $\mathfrak{X}$ contains all groups $G$, for which $kG$ is weakly Gorenstein regular. In particular, $\mathfrak{X}$ contains all groups $G$, for which $kG$ is Gorenstein regular. Hence, the following result generalizes [29, Theorem 2.28].
@@
Let $\mathfrak{X}_{0}$ be a subgroup-closed subclass of $\mathfrak{X}$, such that for any $\mathfrak{X}_{0}$-group $H$ the Gorenstein projective cotorsion pair $\left({\tt GProj}(kH),{\tt GProj}(kH)^{\perp}\right)$ is cogenerated by a set ${\tt S}_{H}$. Consider the closure $\overline{\mathfrak{X}_{0}}$ of $\mathfrak{X}_{0}$ under the operation ${\scriptstyle{{\bf LH}}}\cup\Phi\cup\Phi_{flat}$, as in §$\S$1.IV. Then, for any $\overline{\mathfrak{X}_{0}}$-group $G$ the Gorenstein projective cotorsion pair $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is cogenerated by the set ${\tt T}_{G}=\bigcup\left\{\mbox{ind}_{H}^{G}{\tt S}_{H}:H\subseteq G,H\in\mathfrak{X}_{0}\right\}$.
@@
First of all, we assume that $G\in\mathfrak{X}_{0,0}=\mathfrak{X}_{0}$. Since $\mbox{ind}_{H}^{G}{\tt S}_{H}\subseteq\mbox{ind}_{H}^{G}{\tt GProj}(kH)\subseteq{\tt GProj}(kG)$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$, we have ${\tt T}_{G}\subseteq{\tt GProj}(kG)$. As $G$ is an $\mathfrak{X}_{0}$-subgroup of itself, it is clear that ${\tt S}_{G}\subseteq{\tt T}_{G}$. The inclusions ${\tt S}_{G}\subseteq{\tt T}_{G}\subseteq{\tt GProj}(kG)$ imply that ${\tt GProj}(kG)^{\perp}\subseteq{\tt T}_{G}^{\perp}\subseteq{\tt S}_{G}^{\perp}$. Since ${\tt GProj}(kG)^{\perp}={\tt S}_{G}^{\perp}$, we conclude that ${\tt GProj}(kG)^{\perp}={\tt T}_{G}^{\perp}$, as needed.
@@
for such a subgroup $F$ and all $n>0$. It follows that $\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N^{i}\right)=0$ for all $i$ and all $n>0$. Then, using dimension shifting, we conclude that $\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N\right)=0$ for all $n>d$. Since this holds for any $N\in{\tt GInj}(kH)$, we can easily conclude that the kH-module $\mbox{res}_{H}^{G}M$ is actually contained in ${}^{\perp}{\tt GInj}(kH)$. Indeed, the Gorenstein injective $kH$-module $N$ is the $d$-th cosyzygy in an injective resolution of another Gorenstein injective $kH$-module $N^{\prime}$ (this is property (2’) in Section 2) and hence
@@
(ii) We shall now assume that $\mathfrak{Y}_{0}$ is subgroup-closed and consider an ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-subgroup $H$ of $G$. We proceed by induction on the cardinality $\kappa$ of $H$. If $\kappa$ is countable, then the ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-group $H$ is actually contained in ${\scriptstyle{{\bf H}}}\mathfrak{Y}_{0}$ and we are done by (i) above. If $\kappa$ is uncountable, then we may express $H$ as a continuous ascending union of subgroups $(H_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Since the class ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$ is subgroup-closed, $H_{\lambda}$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-group; hence, our induction hypothesis implies that $\mbox{res}^{G}_{H_{\lambda}}M\in{{}^{\perp}}{\tt GInj}(kH_{\lambda})$ for all $\lambda$. We now fix $N\in{\tt GInj}(kH)$ and note that $\mbox{res}^{H}_{H_{\lambda}}N\in{\tt GInj}(kH_{\lambda})$ for all $\lambda$; this follows since $H_{\lambda}\in{\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{Y}$, in view of property (4’) in Section 2. Letting $M_{\lambda}=\mbox{ind}_{H_{\lambda}}^{H}\mbox{res}^{G}_{H_{\lambda}}M$, we note that
$\mbox{Ext}^{n}_{kH}\!\left(M_{\lambda},N\right)=\mbox{Ext}^{n}_{kH_{\lambda}}\!\left(\mbox{res}^{G}_{H_{\lambda}}M,\mbox{res}_{H_{\lambda}}^{H}N\right)=0$
@@
Here, $T$ is equal to the continuous ascending union of its submodules $(T_{\lambda})_{\lambda<\kappa}$. Since $T_{\lambda+1}/T_{\lambda}$ can be identified with the kernel of the (surjective) structure map $M_{\lambda}\longrightarrow M_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kH}\!\left(T_{\lambda+1}/T_{\lambda},N\right)=0$ for all $\lambda$ and $n>0$. Therefore, Eklof’s lemma [24, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kH}(T,N)=0$ for all $n>0$. Then, the short exact sequence (2) above implies that $\mbox{Ext}^{n}_{kH}\!\left(\mbox{res}_{H}^{G}M,N\right)\!=0$ for all $n>1$. Since this holds for any $N\in{\tt GInj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${{}^{\perp}}{\tt GInj}(kH)$. Indeed, the Gorenstein injective $kH$-module $N$ is the first cosyzygy in an injective resolution of another Gorenstein injective $kH$-module $N^{\prime}$ (this is property (2’) in Section 2) and hence
@@
This completes the proof of (ii).

### v2

@@ 2. A convenient class of groups for studying Gorenstein modules
We fix a commutative ring $k$ and let $\mathfrak{X}=\mathfrak{X}(k)$ be the class consisting of those groups $G$, for which ${\tt PGF}(kG)={\tt GProj}(kG)={\mathcal{P}}(kG)$. In other words, $G$ is an $\mathfrak{X}$-group if all cokernels of acyclic complexes of projective $kG$-modules are projectively coresolved Gorenstein flat. We also consider the class $\mathfrak{X}^{\prime}=\mathfrak{X}^{\prime}(k)$ consisting of those groups $G$, for which ${\tt GFlat}(kG)={\mathcal{F}}(kG)$, and the class $\mathfrak{Y}=\mathfrak{Y}(k)$ consisting of those groups $G$, for which ${\tt GInj}(kG)={\mathcal{I}}(kG)$. In other words, $G$ is an $\mathfrak{X}^{\prime}$-group (resp. a $\mathfrak{Y}$-group) if all cokernels (resp. kernels) of acyclic complexes of flat (resp. injective) $kG$-modules are necessarily Gorenstein flat (resp. Gorenstein injective). These classes have been studied in [23], where the following result is proved.
@@
(i) $\mathfrak{X}=\mathfrak{X}^{\prime}\supseteq\mathfrak{Y}$; cf. [23, Corollary 2.3].
(ii) $\mathfrak{X}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$, $\Phi$ and $\Phi_{flat}$; cf. [23, Theorem 3.3]
(iii) $\mathfrak{Y}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi_{inj}$; cf. [23, Theorem 4.1].
For later use, we record the following properties of Gorenstein projective modules over a group algebra. The first three of these properties are valid over all groups, whereas the latter three are valid under additional assumptions:
@@
(5) If $G$ is an $\mathfrak{X}$-group, then $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is a hereditary cotorsion pair which is cogenerated by a set; in particular, it is complete.
@@
Proof. Property (1) is precisely [22, Lemma 2.6(i)], whereas (2) and (3) are immediate consequences of the definition of Gorenstein projectivity; see also [29, Theorem 2.20]. Property (4) follows since restriction maps $\mathcal{P}(kG)$ into $\mathcal{P}(kH)={\tt GProj}(kH)$, whereas (5) and (6) follow since ${\tt GProj}(kG)={\tt PGF}(kG)$, in view of [37, Theorems 4.4 and 4.9]. $\Box$
@@
(5’) The pair $\left({}^{\perp}{\tt GInj}(kG),{\tt GInj}(kG)\right)$ is a hereditary cotorsion pair which is cogenerated by a set; in particular, it is complete.
Proof. Property (1’) follows by dualizing the argument in [22, Lemma 2.6(i)], whereas (2) is an immediate consequence of the definition of Gorenstein injective modules. Properties (3’) and (5’) follow from [37, Theorem 5.6] and [29, Theorem 2.22]. Finally, property (4’) follows since restriction maps $\mathcal{I}(kG)$ into $\mathcal{I}(kH)={\tt GInj}(kH)$. $\Box$
@@ 3. $\mbox{Ext}^{1}$-orthogonality and group class operations
Let $\mathfrak{X}_{0}$ be a subclass of $\mathfrak{X}$. We consider a group $G$ and a $kG$-module $M$, which is such that $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
@@
(iv) We have $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\Phi_{flat}\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
@@
for such a subgroup $F$ and all $n>0$. It follows that $\mbox{Ext}_{kH}^{n}\!\left(N_{i},\mbox{res}_{H}^{G}M\right)\!=0$ for all $i=0,1,\ldots,d$ and all $n>0$. Using dimension shifting, we conclude that $\mbox{Ext}_{kH}^{n}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>d$. Since this holds for any $N\in{\tt GProj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${\tt GProj}(kH)^{\perp}$. Indeed, the Gorenstein projective $kH$-module $N$ is the $d$-th syzygy in a projective resolution of another Gorenstein projective $kH$-module $N^{\prime}$ (this is property (2) in Section 2) and hence
@@
Here, $K$ is equal to the continuous ascending union of its submodules $(K_{\lambda})_{\lambda<\kappa}$. Since $K_{\lambda+1}/K_{\lambda}$ can be identified with the kernel of the (surjective) structure map $N_{\lambda}\longrightarrow N_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kH}(K_{\lambda+1}/K_{\lambda},\mbox{res}_{H}^{G}M)=0$ for all $\lambda$ and $n>0$. Therefore, Eklof’s lemma [25, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kH}\!\left(K,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>0$. Then, the short exact sequence (1) above implies that $\mbox{Ext}^{n}_{kH}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>1$. Since this holds for any $N\in{\tt GProj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${\tt GProj}(kH)^{\perp}$. Indeed, the Gorenstein projective $kH$-module $N$ is the first syzygy in a projective resolution of another Gorenstein projective $kH$-module $N^{\prime}$ (this is property (2) in Section 2) and hence
@@
We shall prove in the next Section (Proposition 4.3) that $\mathfrak{X}$ contains all groups $G$ for which $kG$ is weakly Gorenstein regular; hence, it contains all groups $G$ for which $kG$ is Gorenstein regular. Therefore, the following result generalizes [31, Theorem 2.28]: For any fixed subgroup-closed subclass $\mathfrak{X}_{0}$ of $\mathfrak{X}$ and any group $G$ which is obtained by iterating the operation ${\scriptstyle{{\bf LH}}}\cup\Phi\cup\Phi_{flat}$ on that subclass, it provides a cogenerating set for the Gorenstein projective cotorsion pair of $kG$, that is controlled by the $\mathfrak{X}_{0}$-subgroups of $G$.
@@
Let $\mathfrak{X}_{0}$ be a subgroup-closed subclass of $\mathfrak{X}$ and assume that for any group $H\in\mathfrak{X}_{0}$ the Gorenstein projective cotorsion pair $\left({\tt GProj}(kH),{\tt GProj}(kH)^{\perp}\right)$ is cogenerated by a set ${\tt S}_{H}$. We also consider the closure $\overline{\mathfrak{X}_{0}}$ of $\mathfrak{X}_{0}$ under the operation ${\scriptstyle{{\bf LH}}}\cup\Phi\cup\Phi_{flat}$, as in §$\S$1.IV. Then, for any $\overline{\mathfrak{X}_{0}}$-group $G$ the Gorenstein projective cotorsion pair $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is cogenerated by the set ${\tt T}_{G}=\bigcup\left\{\mbox{ind}_{H}^{G}{\tt S}_{H}:H\subseteq G,H\in\mathfrak{X}_{0}\right\}$.
@@
First of all, we assume that $G\in\mathfrak{X}_{0,0}=\mathfrak{X}_{0}$. Since $\mbox{ind}_{H}^{G}{\tt S}_{H}\subseteq\mbox{ind}_{H}^{G}{\tt GProj}(kH)\subseteq{\tt GProj}(kG)$ for any ($\mathfrak{X}_{0}$-)subgroup $H\subseteq G$, we have ${\tt T}_{G}\subseteq{\tt GProj}(kG)$. As $G$ is an $\mathfrak{X}_{0}$-subgroup of itself, it is clear that ${\tt S}_{G}\subseteq{\tt T}_{G}$. The inclusions ${\tt S}_{G}\subseteq{\tt T}_{G}\subseteq{\tt GProj}(kG)$ imply that ${\tt GProj}(kG)^{\perp}\subseteq{\tt T}_{G}^{\perp}\subseteq{\tt S}_{G}^{\perp}$. Since ${\tt GProj}(kG)^{\perp}={\tt S}_{G}^{\perp}$, we conclude that ${\tt GProj}(kG)^{\perp}={\tt T}_{G}^{\perp}$, as needed.
@@
for such a subgroup $F$ and all $n>0$. It follows that $\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N^{i}\right)=0$ for all $i$ and all $n>0$. Using dimension shifting, we conclude that $\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N\right)=0$ if $n>d$. Since this holds for any $N\in{\tt GInj}(kH)$, we can easily conclude that the kH-module $\mbox{res}_{H}^{G}M$ is actually contained in ${}^{\perp}{\tt GInj}(kH)$. Indeed, the Gorenstein injective $kH$-module $N$ is the $d$-th cosyzygy in an injective resolution of another Gorenstein injective $kH$-module $N^{\prime}$ (this is property (2’) in Section 2) and hence
@@
(ii) We shall now assume that $\mathfrak{Y}_{0}$ is subgroup-closed and consider an ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-subgroup $H$ of $G$. We proceed by induction on the cardinality $\kappa$ of $H$. If $\kappa$ is countable, then the ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-group $H$ is actually contained in ${\scriptstyle{{\bf H}}}\mathfrak{Y}_{0}$ and we are done by (i) above. If $\kappa$ is uncountable, then we may express $H$ as a continuous ascending union of subgroups $(H_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Since the class ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$ is subgroup-closed, $H_{\lambda}$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-group; our induction hypothesis implies then that $\mbox{res}^{G}_{H_{\lambda}}M\in{{}^{\perp}}{\tt GInj}(kH_{\lambda})$ for all $\lambda$. We now fix $N\in{\tt GInj}(kH)$ and note that $\mbox{res}^{H}_{H_{\lambda}}N\in{\tt GInj}(kH_{\lambda})$ for all $\lambda$; this follows since $H_{\lambda}\in{\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{Y}$, in view of property (4’) in Section 2. Letting $M_{\lambda}=\mbox{ind}_{H_{\lambda}}^{H}\mbox{res}^{G}_{H_{\lambda}}M$, we note that
(2) $\mbox{Ext}^{n}_{kH}\!\left(M_{\lambda},N\right)=\mbox{Ext}^{n}_{kH_{\lambda}}\!\left(\mbox{res}^{G}_{H_{\lambda}}M,\mbox{res}_{H_{\lambda}}^{H}N\right)=0$
@@
Here, $T$ is equal to the continuous ascending union of its submodules $(T_{\lambda})_{\lambda<\kappa}$. Since $T_{\lambda+1}/T_{\lambda}$ can be identified with the kernel of the (surjective) structure map $M_{\lambda}\longrightarrow M_{\lambda+1}$, we conclude from (2) that $\mbox{Ext}^{n}_{kH}\!\left(T_{\lambda+1}/T_{\lambda},N\right)=0$ for all $\lambda$ and $n>0$. Then, Eklof’s lemma [25, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kH}(T,N)=0$ for all $n>0$ and the short exact sequence (3) above implies that $\mbox{Ext}^{n}_{kH}\!\left(\mbox{res}_{H}^{G}M,N\right)\!=0$ for all $n>1$. Since this holds for any $N\in{\tt GInj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${{}^{\perp}}{\tt GInj}(kH)$. Indeed, the Gorenstein injective $kH$-module $N$ is the first cosyzygy in an injective resolution of another Gorenstein injective $kH$-module $N^{\prime}$ (this is property (2’) in Section 2) and hence
@@
This completes the proof of (ii). [^2] One could replace the variation of Benson’s lemma [8, Lemma 5.6] in the argument above and use instead the closure of the left orthogonal class ${{}^{\perp}}{\tt GInj}(kG)$ under colimits (cf. [28, Proposition 3.1]), in order to conclude from (2) that the colimit $\mbox{res}_{H}^{G}M$ of the direct system $(M_{\lambda})_{\lambda<\kappa}$ is also contained in ${{}^{\perp}}{\tt GInj}(kG)$.